OpenStudy (anonymous):

Kyle had two large sheets of paper .He cut each piece of paper in half, then he cut the smaller pieces of paper in half two more times. Write this as a power!

OpenStudy (anonymous):

help someone..................

OpenStudy (anonymous):

It's half of a half of a half of a whole. Thus: \[\left(\dfrac{1}{2}\right) \left(\dfrac{1}{2}\right) \left(\dfrac{1}{2}\right)1=\dfrac{1}{2^3}=2^{-3}\]For more solutions to textbook problems, be sure to check out http://www.slader.com/s/eWFrZXlnbGVl . I submit a lot of content there as well.

OpenStudy (anonymous):

Whoops sorry, he had two sheets of paper: thus we have a half of a half of a half of two wholes: \[\left(\dfrac 12\right)\left(\dfrac 12\right)\left(\dfrac 12\right)2 = \dfrac{2}{2^3}=\dfrac{1}{2^2}=2^{-2}\]

OpenStudy (anonymous):

thanks alot really helped!